OpenStudy (anonymous):

multiply x^2+8x+15/x-4 times x^2-16/2x+6

OpenStudy (lgbasallote):

ahh i like this one >:)) let us factor each one first x^2 + 8x + 5 -> x- 4 -> x^2 - 16 -> 2x + 6 -> can you factor those?

OpenStudy (anonymous):

does x^2 and x have factors? i know that sounds dumb but i just want to make sure..

OpenStudy (lgbasallote):

yep.. i.e x^2 + x = x(x + 1)

OpenStudy (lgbasallote):

but where are you going to use that?

OpenStudy (anonymous):

I dont thik I need to use it

OpenStudy (lgbasallote):

ahh yes you wont hehe...so how will you factor the 4 stated expressions?

OpenStudy (anonymous):

I might need some help with that

OpenStudy (lgbasallote):

let's start with 2x + 6 since it's the easiest...do you see any factors?

OpenStudy (anonymous):

2

OpenStudy (lgbasallote):

good! so if you factor out 2..what will you get?

OpenStudy (anonymous):

uhm.. 1?

OpenStudy (lgbasallote):

this is how you'll now \(\LARGE 2(\frac{2x}{2} + \frac{6}[2}\) you divide each term by 2

OpenStudy (lgbasallote):

\(2(\frac{2x}{2} + \frac{6}{2})\)

OpenStudy (lgbasallote):

there we go...so what's 2x/2? and 6/2?

OpenStudy (anonymous):

x and3

OpenStudy (lgbasallote):

so (2x+6) = 2(x+3) got it?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

let's move on to the next one... \(x^2 - 16\) Hint: \(\large (a+b)(a-b) = a^2 - b^2 \leftarrow \text{difference of two squares}\)

OpenStudy (anonymous):

im not sure

OpenStudy (lgbasallote):

okay...in x^2 - 16 what is your "a^2"?

OpenStudy (anonymous):

im sorry im still not sure

OpenStudy (lgbasallote):

hmm you understand that x^2 - 16 is in the form of a^2 - b^2 right?

OpenStudy (anonymous):

i didnt know that

OpenStudy (lgbasallote):

wll x^2 - 16 is in that form...you want to change it to (a+b)(a-b) but to do that you need to find a and b...following?

OpenStudy (anonymous):

yes got it

OpenStudy (lgbasallote):

so since x^2 - 16 is in that form we assume that a^2 = x^2 and b^2 = 16 got it? because a^2 is in the left of the minus sign, so is x^2. b^2 is in the right of the minus sign, so is 16. got it?

OpenStudy (anonymous):

yesh

OpenStudy (lgbasallote):

so we know a^2 = x^2 to find a we square root both sides..what's the answer?

OpenStudy (anonymous):

sqare root 16 right?

OpenStudy (lgbasallote):

that's for b...but for a it's square root of x^2...you can simplify sqrt x^2 further..what is it?

OpenStudy (anonymous):

it would just be x right?

OpenStudy (lgbasallote):

right! then b= sqrt 16...what is it simplified further?

OpenStudy (anonymous):

4

OpenStudy (lgbasallote):

so a = x and b = 4 what is (a+b)(a-b) now?

OpenStudy (anonymous):

(x+4)(x-4)

OpenStudy (lgbasallote):

you got it! so let's move on to the next one...the next one is x - 4 can you simplify that further?

OpenStudy (anonymous):

i dont think so

OpenStudy (lgbasallote):

right! so we're down to our last one... x^2 + 8x + 3 do you have any ideas how to factor this out?

OpenStudy (lgbasallote):

oh wait..that's x^2 + 8x + 15 pardon the typo

OpenStudy (anonymous):

so x, 2 and 5?

OpenStudy (lgbasallote):

uhmm howd you get that?

OpenStudy (anonymous):

yeah i dont hink i did that right

OpenStudy (lgbasallote):

well i'll link you to one of my tutorials you may get an idea from there http://openstudy.com/users/lgbasallote#/updates/4f9bd87ee4b000ae9ed11143

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

wow dude -___- i just did 44 minutes trying to teach that :/

OpenStudy (anonymous):

Thanks for showing me and helping me :)

OpenStudy (anonymous):

@ igbasallote

OpenStudy (lgbasallote):

that's the last step anyway..the next steps was what he showed.. and it's @lgbasallote lowercase L

OpenStudy (anonymous):

@Mikieangelo this may help